### v2

@@ 5 Case studies on example models
The most naive method is to construct the sought-after UV limit measure $\mu$ from a fixed reference (Gaussian) measure $\gamma$ on $\mathcal{S}^{\prime}$ as described in Section 1, and from an interaction potential $V:\,\mathcal{S}^{\prime}\to\overline{\mathbb{R}}$, according to $\mu:=\mathrm{e}^{-V}\cdot\gamma$. Here $V$ is some kind of Borel measurable function, associated to the naive potential $\mathcal{V}:\,\mathcal{E}\cap\mathcal{S}^{\prime}\to\overline{\mathbb{R}}$, usually defined by some local expression similar to that of $\mathcal{V}(\varphi)=g\,\int_{\Omega}\varphi^{4}$, with $\Omega$ being some bounded open region of $\mathbb{R}^{N}$ (IR cutoff). The most simple association procedure of $V$ to $\mathcal{V}$ is by smoothing (mollifying), i.e. taking the limit of $\mathcal{V}\circ C_{\eta_{n}}$ in a suitable sense as $(\eta_{n})_{n\in\mathbb{N}}$ is an approximate identity in $\mathcal{S}$. To that end, let us introduce the set of mollifier sequences
@@
Note the identity $\mathrm{e}^{-V}=\mathop{(\gamma)\!\sup}\limits_{(\eta_{n})\in\Delta}\;\limsup\limits_{n\to\infty}\;\mathrm{e}^{-\mathcal{V}\circ C_{\eta_{n}}}$ if $V$ is the greedy extension of $\mathcal{V}$. Therefore, the above is the most economic measurable extension of potentials via smoothing, optimized such that $\mathrm{e}^{-V}$ has the largest possible overlap integral with $\gamma$ on all Borel sets. The following theorem summarizes its basic properties.
@@
Let $\gamma$ be a reference measure, and let $\mathcal{V}:\,\mathcal{E}\cap\mathcal{S}^{\prime}\to\overline{\mathbb{R}}$ be a measurable function bounded from below. Define $V$ to be its greedy extension according to (32). Then $V$ is also a measurable function $\gamma$-essentially bounded from below, thus $\mu:=\mathrm{e}^{-V}\cdot\gamma$ defines a sigma-additive non-negative valued finite measure. The measure $\mu$ is not the zero measure if and only if
$\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathop{(\gamma)\!\sup}\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\,\mathrm{e}^{-\mathcal{V}(\eta_{n}\star\phi)}\,\mathrm{d}\gamma(\phi)\;>\;0.$ (33)
@@
$\sup\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathrm{e}^{-\mathcal{V}(\eta_{n}\star\phi)}\,\mathrm{d}\gamma(\phi)\;>\;0$ (34)
@@
Using the knowledge that $\mathcal{V}$ is bounded from below, the family of functions $\mathrm{e}^{-\mathcal{V}\circ C_{\eta}}$ ($\eta\in\mathcal{S}$) are bounded from above by some $K\geq 0$. That is, they have a $\gamma$-integrable majorant $K$. Taking now an approximate identity $(\eta_{n})\in\Delta$, and applying the Fatou lemma ([34] Theorem 1.28) to the sequence of non-negative valued measurable functions $K-\mathrm{e}^{-\mathcal{V}\circ C_{\eta_{n}}}$, then taking the negative of both sides of the Fatou inequality, and adding $\int_{\phi\in\mathcal{S}^{\prime}}K\,\mathrm{d}\gamma(\phi)$, one infers
$\limsup\limits_{n\to\infty}\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathrm{e}^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi)\;\leq\;\int\limits_{\phi\in\mathcal{S}^{\prime}}\limsup\limits_{n\to\infty}\mathrm{e}^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi).$ (35)
Since $\limsup\limits_{n\to\infty}\mathrm{e}^{-\mathcal{V}\circ C_{\eta_{n}}}\leq\mathop{(\gamma)\!\sup}\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\mathrm{e}^{-\mathcal{V}\circ C_{\eta_{n}}}$ holds $\gamma$-almost everywhere, by applying the monotonity of integration to the right hand side of (35), one has
$\limsup\limits_{n\to\infty}\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathrm{e}^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi)\;\leq\;\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathop{(\gamma)\!\sup}\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\mathrm{e}^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi).$ (36)
@@
$\sup\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathrm{e}^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi)\;\leq\;\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathop{(\gamma)\!\sup}\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\mathrm{e}^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi)$ (37)
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\;=\;\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathrm{e}^{-V(\phi)}\,\mathrm{d}\gamma(\phi)$ (38)
@@
Let $\gamma$ be a reference measure, and let $\mathcal{V}:\,\mathcal{E}\cap\mathcal{S}^{\prime}\to\overline{\mathbb{R}}$ be a measurable function bounded from below. Let $V$ be its greedy extension according to (32), and define the corresponding sigma-additive non-negative valued finite measure $\mu:=\mathrm{e}^{-V}\cdot\gamma$. Then, $\mu$ is not the zero measure if and only if there exists some approximate identity $(\eta_{n})_{n\in\mathbb{N}}$ in $\mathcal{S}$, such that
$\int\limits_{\phi\in\mathcal{S}^{\prime}}\limsup\limits_{n\to\infty}\mathrm{e}^{-\mathcal{V}(\eta_{n}\star\phi)}\,\mathrm{d}\gamma(\phi)\;>\;0$ (39)
@@
$\lim\limits_{n\to\infty}\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathrm{e}^{-\mathcal{V}(\eta_{n}\star\phi)}\,\mathrm{d}\gamma(\phi)\;>\;0$ (40)
@@
The above necessary and sufficient condition (39) for the nonvanishing of $\mathrm{e}^{-V}\cdot\gamma$ is typically hard to evaluate, given some $\mathcal{V}$. The sufficient condition (40), however, is in principle evaluable for concrete QFT models, by mapping $\mathcal{S}^{\prime}$, $\gamma$ and the operations defining $\mathcal{V}$ to sequence spaces. It is still an unresolved interesting question, that along which additional properties on $\mathcal{V}$ the condition (40) would be also necessary. If such properties could be uncovered, it would provide a rather powerful tool for checking the renormalizability of concrete QFT models. For potentials having finite UV limit of their expectation value, i.e. in case
@@
holds for some approximate identity $(\eta_{n})_{n\in\mathbb{N}}$ in $\mathcal{S}$, Jensen’s inequality ([34] Theorem 3.3) implies that the sufficient condition (40) is satisfied. Evaluating this, it follows that for the IR cutoff $\varphi^{4}$ model over a 0+1 dimensional spacetime, the greedy extension defines a valid interacting Feynman measure. [^4] It is of course well known in the literature that this model has a Feynman measure over $\mathbb{R}^{1}$ spacetime. That is simply because a free Klein–Gordon Gaussian measure on fields over $\mathbb{R}^{1}$ is supported on the continuous functions, see Remark 37 (ii) in A. The condition (41) is also satisfied for IR cutoff potentials in arbitrary dimensional spacetimes whenever the potential density is bounded both from below and above. Such an example is the potential $\mathcal{V}(\varphi)=g\,\int_{\Omega}\frac{1}{1+l^{4}\,\varphi^{4}}\,\varphi^{4}$ (“Weidling” potential), or a sine–Gordon potential. Therefore, with the greedy extension, these define their Feynman measures in a relatively straightforward way. The greedy extension preserves lower bound, and therefore is particularly attractive for extension of regularized potentials which are themselves bounded from below, as it respects the constraint by Theorem 25.
@@
$W\left(\phi\right)=\begin{cases}\mathcal{V}\left(\phi\right)&\text{if }\phi\in\mathcal{E}\cap\mathcal{S}^{\prime}\\ +\infty&\text{otherwise,}\end{cases}$ (42)
@@
A straightforward consequence is that all sublevel sets $\overset{-1}{\underline{W}}([-\infty,t])$ for $t\in\mathbb{R}$ are weakly closed in $\mathcal{S}^{\prime}$. Hence, $\underline{W}$ is automatically Borel measurable.
A benefit of this construction is that it is always smaller than any (meaningful) limit of a sequence $\mathcal{V}\circ C_{\eta_{n}}$ ($n\in\mathbb{N}$), in particular it is smaller than the greedy extension. Consequently, $\mathrm{e}^{-\underline{W}}$ has a greater chance of not being zero $\gamma$-almost everywhere, in contrast to a corresponding limit of $\mathcal{V}\circ C_{\eta_{n}}$. If $\mathcal{V}$ was bounded from below, then so is its lower semicontinuous envelope $\underline{W}$, and thus $\mathrm{e}^{-\underline{W}}\cdot\gamma$ defines a sigma-additive non-negative valued finite measure. Similar to the greedy extension, this product is nonzero whenever the overlap integral of $\mathrm{e}^{-\underline{W}}$ and $\gamma$ is nonzero.
It is apparent that the construction of UV limit potentials in general is very difficult. In that spirit, below we collect some frequent pathology properties of a potential $\mathcal{V}$, such that even taking its lower semicontinuous envelope $\underline{W}$ as above does not help to define a nonzero measure $\mathrm{e}^{-\underline{W}}\cdot\gamma$.
@@
$W:\,Y\to\overline{\mathbb{R}},\qquad\phi\mapsto\begin{cases}\mathcal{V}(\iota^{-1}(\phi))&\text{if }\phi\in\iota[X]\\ +\infty&\text{else}\end{cases}$ (46)
@@
Here $\widehat{(\cdot)}$ denotes the canonical injection of a vector into the continuous bidual space. Note that $j^{**}$ is equal to the weak-*-to-weak continuous extension of $j$ to the bidual space $\pi[X]^{\prime\prime}$, hence $\pi(\varphi_{\beta})=j\pi(\varphi_{\beta})=j^{**}\widehat{\pi(\varphi_{\beta}^{\prime})}\to j^{**}z$ weakly in $Y$. Since $\iota\varphi_{\beta}^{\prime}\to\phi$ weakly in $Y$, it follows that $(\iota-\pi)\varphi_{\beta}^{\prime}\to\phi-j^{**}z$ weakly in Y, thus
@@
$W:\,\mathcal{S}^{\prime}\to\overline{\mathbb{R}},\qquad\phi\mapsto\begin{cases}\mathcal{V}(\varphi)&\text{when }\phi=\iota\varphi\text{ for some }\varphi\in\mathcal{S}\\ +\infty&\text{otherwise}\end{cases}$ (48)
@@
Observe, that the above no-go theorem is avoided by the aforementioned “Weidling” potential, as well as the competing Higgs potentials of the form $\mathcal{V}(\varphi,\psi)=g\,\int_{\Omega}(\varphi^{2}-\psi^{2})^{2}$, since their corresponding characterizing sets $L$ from (45) will contain finite values (in particular, the zero). That is, the pertinent potentials are not coercive with respect to the $L^{4}$ norm. The IR cutoff $\varphi^{4}$ potential over 1 dimensional spacetime is of course coercive with respect to $L^{4}$, but the no-go theorem is avoided because $L^{4}(\Omega)\subset\mathcal{S}^{\prime}$ is not zero $\gamma$-measure subspace in that setting.
@@
$\mathbb{E}_{\gamma}[\cdot]$ denoting the $\gamma$-expectation value. Over $\mathbb{R}^{2}$ spacetimes, it is known that $V_{n.o.,\eta}$ converges to some $V_{n.o.}$ in the $L^{2}_{\gamma}$ norm as $\eta\to\delta$ (as approximate identity), and that the corresponding $\mathrm{e}^{-V_{n.o.}}$ is $\gamma$-integrable (see a review in [38]). Therefore, this $\mathrm{e}^{-V_{n.o.}}\cdot\gamma$ defines a Feynman measure for the IR cutoff $\varphi^{4}$ model over 2 dimensional spacetime. One should note, however, that the family of regularized measures $\mathrm{e}^{-V_{n.o.,\eta}}\cdot\gamma$, used in the above constructive limiting procedure, itself is not a Wilsonian RG flow: they are not marginals to one another with respect to some coarse-graining and field rescaling. Of course, after the limiting measure $\mathrm{e}^{-V_{n.o.}}\cdot\gamma$ is constructed, it induces a Wilsonian RG flow, and it is an interesting open question what relation it admits with respect to the family of measures $\mathrm{e}^{-V_{n.o.,\eta}}\cdot\gamma$ ($\eta\in\mathcal{S}$).
@@
was considered, with the real valued continuous nonlinear functionals $a,b:\,\mathcal{S}\to\mathbb{R}$ of the regulators being not a priori fixed. These are sometimes called running couplings and their corresponding terms as counterterms. Clearly, normal ordering is a special case of this ansatz. The unkown functionals $a,b$ are fixed by additional conditions on certain observables, sometimes referred to as renormalization conditions. In [21, 22] it was shown that the $\varphi^{4}$ potential of the above kind over $\mathbb{R}^{3}$ spacetime admits a weak limit measure $\mu$ of the family of regularized measures $\mathrm{e}^{-V_{a(\eta),b(\eta),\eta}}\cdot\gamma$ as $\eta\to\delta$ (as approximate identity), whenever the functionals $a,b$ are suitably chosen. Although, the explicit form of the suitable functionals $a,b$ for this model are not yet known, it was shown that the limiting measure $\mu$ cannot be absolutely continuous with respect to the reference measure $\gamma$ one started from. Therefore, UV limit of the potential family $V_{a(\eta),b(\eta),\eta}$ ($\eta\in\mathcal{S}$) does not exist. Similarly to Example 34, the pertinent family of measures used in the limiting procedure is not a Wilsonian RG flow itself, but given the limiting measure $\mu$ it induces a Wilsonian RG flow.
We note in passing that similar construction to the above was worked out in [39] for $\varphi^{4}$ potential over $\mathbb{R}^{4}$, but with a lattice spacetime regularization, and therefore it is not discussed in the present paper, as we concentrate on continuum spacetimes.
@@
One could also consider an even more general limiting procedure such that both the regularized potential as well as the reference measure depends on the regularization $\eta\in\mathcal{S}$. This scenario can happen, for instance, if the reference measure $\gamma_{d(\eta),m(\eta)}$ is a Gaussian measure subordinate to the Klein–Gordon operator $(-d^{2}(\eta)\,\Delta+m^{2}(\eta))$, and where the $d$ parameter as well as the mass parameter $m$ is taken to be a not yet specified $\eta$-dependent coefficient, similar to that of the running couplings in the potential of Example 35. The regularized measure $\mathrm{e}^{-V_{\eta}}\cdot\gamma_{d(\eta),m(\eta)}$ then may have some weak limit measure $\mu$ as $\eta\to\delta$ (as approximate identity), whenever the running couplings in $V_{\eta}$, $d(\eta)$ and the running mass $m(\eta)$ are suitably chosen. For this case, we are not aware of a mathematically worked out example in the literature. Again, such limiting procedure itself is non-Wilsonian, but the limit induces a Wilsonian RG flow.
In conclusion, based on Corollary 15, Corollary 21 and Theorem 25, one can state that if particularly the Wilsonian RG prescription is important for the description of interactions appearing in Nature, and the interacting measure is to be described by a potential with respect to a Gaussian reference flow induced by a Klein–Gordon operator, then in 4 dimensional spacetimes certain interactions, such as the $\varphi^{4}$ are disfavoured, whereas other interactions such as the “Weidling” potential or the competing Higgs model are favoured.

### v3

@@ 5 Case studies on example models
The most naive method is to construct the sought-after UV limit measure $\mu$ from a fixed reference (Gaussian) measure $\gamma$ on $\mathcal{S}^{\prime}$ as described in Section 1, and from an interaction potential $V:\,\mathcal{S}^{\prime}\to\overline{\mathbb{R}}$, according to $\mu:=\rme^{-V}\cdot\gamma$. Here $V$ is some kind of Borel measurable function, associated to the naive potential $\mathcal{V}:\,\mathcal{E}\cap\mathcal{S}^{\prime}\to\overline{\mathbb{R}}$, usually defined by some local expression similar to that of $\mathcal{V}(\varphi)=g\,\int_{\Omega}\varphi^{4}$, with $\Omega$ being some bounded open region of $\mathbb{R}^{N}$ (IR cutoff). The most simple association procedure of $V$ to $\mathcal{V}$ is by smoothing (mollifying), i.e. taking the limit of $\mathcal{V}\circ C_{\eta_{n}}$ in a suitable sense as $(\eta_{n})_{n\in\mathbb{N}}$ is an approximate identity in $\mathcal{S}$. To that end, let us introduce the set of mollifier sequences
@@
Note the identity $\rme^{-V}=\mathop{(\gamma)\!\sup}\limits_{(\eta_{n})\in\Delta}\;\limsup\limits_{n\to\infty}\;\rme^{-\mathcal{V}\circ C_{\eta_{n}}}$ if $V$ is the greedy extension of $\mathcal{V}$. Therefore, the above is the most economic measurable extension of potentials via smoothing, optimized such that $\rme^{-V}$ has the largest possible overlap integral with $\gamma$ on all Borel sets. The following theorem summarizes its basic properties.
@@
Let $\gamma$ be a reference measure, and let $\mathcal{V}:\,\mathcal{E}\cap\mathcal{S}^{\prime}\to\overline{\mathbb{R}}$ be a measurable function bounded from below. Define $V$ to be its greedy extension according to (34). Then $V$ is also a measurable function $\gamma$-essentially bounded from below, thus $\mu:=\rme^{-V}\cdot\gamma$ defines a sigma-additive non-negative valued finite measure. The measure $\mu$ is not the zero measure if and only if
$\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathop{(\gamma)\!\sup}\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\,\rme^{-\mathcal{V}(\eta_{n}\star\phi)}\,\mathrm{d}\gamma(\phi)\;>\;0.$ (35)
@@
$\sup\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\int\limits_{\phi\in\mathcal{S}^{\prime}}\rme^{-\mathcal{V}(\eta_{n}\star\phi)}\,\mathrm{d}\gamma(\phi)\;>\;0$ (36)
@@
Using the knowledge that $\mathcal{V}$ is bounded from below, the family of functions $\rme^{-\mathcal{V}\circ C_{\eta}}$ ($\eta\in\mathcal{S}$) are bounded from above by some $K\geq 0$. That is, they have a $\gamma$-integrable majorant $K$. Taking now an approximate identity $(\eta_{n})\in\Delta$, and applying the Fatou lemma ([39] Theorem 1.28) to the sequence of non-negative valued measurable functions $K-\rme^{-\mathcal{V}\circ C_{\eta_{n}}}$, then taking the negative of both sides of the Fatou inequality, and adding $\int_{\phi\in\mathcal{S}^{\prime}}K\,\mathrm{d}\gamma(\phi)$, one infers
$\limsup\limits_{n\to\infty}\int\limits_{\phi\in\mathcal{S}^{\prime}}\rme^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi)\;\leq\;\int\limits_{\phi\in\mathcal{S}^{\prime}}\limsup\limits_{n\to\infty}\rme^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi).$ (37)
Since $\limsup\limits_{n\to\infty}\rme^{-\mathcal{V}\circ C_{\eta_{n}}}\leq\mathop{(\gamma)\!\sup}\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\rme^{-\mathcal{V}\circ C_{\eta_{n}}}$ holds $\gamma$-almost everywhere, by applying the monotonity of integration to the right hand side of (37), one has
$\limsup\limits_{n\to\infty}\int\limits_{\phi\in\mathcal{S}^{\prime}}\rme^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi)\;\leq\;\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathop{(\gamma)\!\sup}\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\rme^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi).$ (38)
@@
$\sup\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\int\limits_{\phi\in\mathcal{S}^{\prime}}\rme^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi)\;\leq\;\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathop{(\gamma)\!\sup}\limits_{(\eta_{n})\in\Delta}\limsup\limits_{n\to\infty}\rme^{-\mathcal{V}(C_{\eta_{n}}\,\phi)}\,\mathrm{d}\gamma(\phi)$ (39)
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\;=\;\int\limits_{\phi\in\mathcal{S}^{\prime}}\rme^{-V(\phi)}\,\mathrm{d}\gamma(\phi)$ (40)
@@
Let $\gamma$ be a reference measure, and let $\mathcal{V}:\,\mathcal{E}\cap\mathcal{S}^{\prime}\to\overline{\mathbb{R}}$ be a measurable function bounded from below. Let $V$ be its greedy extension according to (34), and define the corresponding sigma-additive non-negative valued finite measure $\mu:=\rme^{-V}\cdot\gamma$. Then, $\mu$ is not the zero measure if and only if there exists some approximate identity $(\eta_{n})_{n\in\mathbb{N}}$ in $\mathcal{S}$, such that
$\int\limits_{\phi\in\mathcal{S}^{\prime}}\limsup\limits_{n\to\infty}\rme^{-\mathcal{V}(\eta_{n}\star\phi)}\,\mathrm{d}\gamma(\phi)\;>\;0$ (41)
@@
$\lim\limits_{n\to\infty}\int\limits_{\phi\in\mathcal{S}^{\prime}}\rme^{-\mathcal{V}(\eta_{n}\star\phi)}\,\mathrm{d}\gamma(\phi)\;>\;0$ (42)
@@
The above necessary and sufficient condition (41) for the nonvanishing of $\rme^{-V}\cdot\gamma$ is typically hard to evaluate, given some $\mathcal{V}$. The sufficient condition (42), however, is in principle evaluable for concrete QFT models, by mapping $\mathcal{S}^{\prime}$, $\gamma$ and the operations defining $\mathcal{V}$ to sequence spaces. It is still an unresolved interesting question, that along which additional properties on $\mathcal{V}$ the condition (42) would be also necessary. If such properties could be uncovered, it would provide a rather powerful tool for checking the renormalizability of concrete QFT models. For potentials having finite UV limit of their expectation value, i.e. in case
@@
holds for some approximate identity $(\eta_{n})_{n\in\mathbb{N}}$ in $\mathcal{S}$, Jensen’s inequality ([39] Theorem 3.3) implies that the sufficient condition (42) is satisfied. Evaluating this, it follows that for the IR cutoff $\varphi^{4}$ model over a 0+1 dimensional spacetime, the greedy extension defines a valid interacting Feynman measure. [^5] It is of course well known in the literature that this model has a Feynman measure over $\mathbb{R}^{1}$ spacetime. That is simply because a free Klein–Gordon Gaussian measure on fields over $\mathbb{R}^{1}$ is supported on the continuous functions, see Remark 37 (ii) in A. The condition (43) is also satisfied for IR cutoff potentials in arbitrary dimensional spacetimes whenever the potential density is bounded both from below and above. Such an example is the potential $\mathcal{V}(\varphi)=g\,\int_{\Omega}\frac{1}{1+l^{4}\,\varphi^{4}}\,\varphi^{4}$ (“basin-shaped” potential), or a sine–Gordon potential. [^6] A very recent result on the triviality of Feynman measures subordinate to certain bounded potentials is proved in [27]. Therefore, with the greedy extension, these define their Feynman measures in a relatively straightforward way. The greedy extension preserves lower bound, and therefore is particularly attractive for extension of regularized potentials which are themselves bounded from below, as it respects the constraint by Theorem 25.
@@
$W\left(\phi\right)=\cases{\mathcal{V}\left(\phi\right)&if $\phi\in\mathcal{E}\cap\mathcal{S}^{\prime}$\\ +\infty&otherwise,}$ (44)
@@
[…]
A benefit of this construction is that it is always smaller than any (meaningful) limit of a sequence $\mathcal{V}\circ C_{\eta_{n}}$ ($n\in\mathbb{N}$), in particular it is smaller than the greedy extension. Consequently, $\rme^{-\underline{W}}$ has a greater chance of not being zero $\gamma$-almost everywhere, in contrast to a corresponding limit of $\mathcal{V}\circ C_{\eta_{n}}$. If $\mathcal{V}$ was bounded from below, then so is its lower semicontinuous envelope $\underline{W}$, and thus $\rme^{-\underline{W}}\cdot\gamma$ defines a sigma-additive non-negative valued finite measure. Similar to the greedy extension, this product is nonzero whenever the overlap integral of $\rme^{-\underline{W}}$ and $\gamma$ is nonzero.
It is apparent that the construction of UV limit potentials in general is very difficult. In that spirit, below we collect some frequent pathology properties of a potential $\mathcal{V}$, such that even taking its lower semicontinuous envelope $\underline{W}$ as above does not help to define a nonzero measure $\rme^{-\underline{W}}\cdot\gamma$.
@@
$W:\,Y\to\overline{\mathbb{R}},\qquad\phi\mapsto\cases{\mathcal{V}(\iota^{-1}(\phi))&if $\phi\in\iota[X]$\\ +\infty&otherwise}$ (48)
@@
Here $\widehat{(\cdot)}$ denotes the canonical injection of a vector into the continuous bidual space. Note that $j^{**}$ is equal to the weak-*-to-weak continuous extension of $j$ to the bidual space $\pi[X]^{\prime\prime}$, hence $\pi(\varphi_{\beta}^{\prime})=j\pi(\varphi_{\beta}^{\prime})=j^{**}\widehat{\pi(\varphi_{\beta}^{\prime})}\to j^{**}z$ weakly in $Y$. Since $\iota\varphi_{\beta}^{\prime}\to\phi$ weakly in $Y$, it follows that $(\iota-\pi)\varphi_{\beta}^{\prime}\to\phi-j^{**}z$ weakly in Y, thus
@@
$W:\,\mathcal{S}^{\prime}\to\overline{\mathbb{R}},\qquad\phi\mapsto\cases{\mathcal{V}(\varphi)&when $\phi=\iota\varphi$ for some $\varphi\in\mathcal{S}$\\ +\infty&otherwise}$ (50)
@@
Observe, that the above no-go theorem is avoided by the aforementioned basin-shaped potential, as well as the competing Higgs potentials of the form $\mathcal{V}(\varphi,\psi)=g\,\int_{\Omega}(\varphi^{2}-\psi^{2})^{2}$, since their corresponding characterizing sets $L$ from (47) will contain finite values (in particular, the zero). That is, the pertinent potentials are not coercive with respect to the $L^{4}$ norm. The IR cutoff $\varphi^{4}$ potential over 1 dimensional spacetime is of course coercive with respect to $L^{4}$, but the no-go theorem is avoided because $L^{4}(\Omega)\subset\mathcal{S}^{\prime}$ is not zero $\gamma$-measure subspace in that setting.
@@
$\mathbb{E}_{\gamma}[\cdot]$ denoting the $\gamma$-expectation value. Over $\mathbb{R}^{2}$ spacetimes, it is known that $V_{n.o.,\eta}$ converges to some $V_{n.o.}$ in the $L^{2}_{\gamma}$ norm as $\eta\to\delta$ (as approximate identity), and that the corresponding $\rme^{-V_{n.o.}}$ is $\gamma$-integrable (see a review in [17]). Therefore, this $\rme^{-V_{n.o.}}\cdot\gamma$ defines a Feynman measure for the IR cutoff $\varphi^{4}$ model over 2 dimensional spacetime. One should note, however, that the family of regularized measures $\rme^{-V_{n.o.,\eta}}\cdot\gamma$, used in the above constructive limiting procedure, itself is not a Wilsonian RG flow: they are not marginals to one another with respect to some coarse-graining and field rescaling (as they violate Theorem 25). Of course, after the limiting measure $\rme^{-V_{n.o.}}\cdot\gamma$ is constructed by any procedure, it then induces a Wilsonian RG flow, and it is an interesting open question what relation it admits with respect to the family of measures $\rme^{-V_{n.o.,\eta}}\cdot\gamma$ ($\eta\in\mathcal{S}$) used in the constructive limiting procedure.
@@
was taken, with the real valued continuous nonlinear functionals $a,b:\,\mathcal{S}\to\mathbb{R}$ of the regulators being not a priori fixed. These are sometimes called running couplings and their corresponding terms as counterterms. Clearly, normal ordering is a special case of this ansatz. The unknown functionals $a,b$ are fixed by additional conditions on certain observables, sometimes referred to as renormalization conditions. In [18, 19] it was shown that the $\varphi^{4}$ potential of the above kind over $\mathbb{R}^{3}$ spacetime admits a weak limit measure $\mu$ of the family of regularized measures $\rme^{-V_{a(\eta),b(\eta),\eta}}\cdot\gamma$ as $\eta\to\delta$ (as approximate identity), whenever the functionals $a,b$ are suitably chosen. Although, the explicit form of the suitable functionals $a,b$ for this model are not yet known, it was shown that the limiting measure $\mu$ cannot be absolutely continuous with respect to the reference measure $\gamma$ one started from. Therefore, UV limit of the potential family $V_{a(\eta),b(\eta),\eta}$ ($\eta\in\mathcal{S}$) does not exist. Similarly to Example 34, the pertinent family of measures used in the limiting procedure is not a Wilsonian RG flow itself (as they violate Corollary 15), but given the limiting measure $\mu$ it induces a Wilsonian RG flow.
We note in passing that a very famous recent construction similar to the above was worked out in [20] for $\varphi^{4}$ potential over $\mathbb{R}^{4}$, but with a lattice spacetime regularization. This case is not discussed in the present paper, as we concentrate on continuum spacetimes.
@@
One could also consider an even more general limiting procedure such that both the regularized potential as well as the reference measure depends on the regularization $\eta\in\mathcal{S}$. This scenario can happen, for instance, if the reference measure $\gamma_{d(\eta),m(\eta)}$ is a Gaussian measure subordinate to the Klein–Gordon operator $(-d^{2}(\eta)\,\Delta+m^{2}(\eta))$, and where the $d$ parameter as well as the mass parameter $m$ is taken to be a not yet specified $\eta$-dependent coefficient, similar to that of the running couplings in the potential of Example 35. The regularized measure $\rme^{-V_{\eta}}\cdot\gamma_{d(\eta),m(\eta)}$ then may have some weak limit measure $\mu$ as $\eta\to\delta$ (as approximate identity), whenever the running couplings in $V_{\eta}$, $d(\eta)$ and the running mass $m(\eta)$ are suitably chosen. For this case, we are not aware of a mathematically worked out example in the literature. Again, such limiting procedure itself is non-Wilsonian, but the limit induces a Wilsonian RG flow.
In conclusion, based on Corollary 15, Corollary 21 and Theorem 25, one can state that if particularly the Wilsonian RG prescription is important for the constructive description of interactions appearing in Nature, and the interacting measure is to be described by a potential with respect to a Gaussian reference flow induced by a Klein–Gordon operator, then in 4 dimensional spacetimes certain interactions, such as the $\varphi^{4}$ are disfavored, whereas other interactions such as the basin-shaped potential, or the $(\varphi^{2}-\psi^{2})^{2}$ type competing Higgs model are favored.